A certain Disk has 400 sectors in a track and its spindle rotates at the speed of 3600 revolutions per minute. If the average seek time and latency time are 7.52 milliseconds and 3.75, respectively, compute the total time it takes to read a particular sector and access a record.


1
Expert's answer
2021-02-15T23:33:01-0500

The total time to read a particular sector and access a record is the sum of the seek time, rotational latency time and transfer time. We already know the seek time and rotational latency time from the condition of the question. Let's find the transfer time. The transfer time is the time it takes for the head to read all of the sector. Since there are on average 400 sectors in a track, we need 1/400th of a revolution of the disk to read the entire sector:


"t=\dfrac{Time\ for\ one\ revolution\ of\ disk}{Number\ of\ sectors\ in\ track},""t=\dfrac{\dfrac{\dfrac{1000\ ms}{1\ s}\cdot\dfrac{60\ s}{1\ min}}{3600\ RPM}}{400}=0.042\ ms."

Therefore, the total time is 7.52 ms + 3.75 ms + 0.042 ms =11.3 ms.